# The Origin Symphony: Probing Baryogenesis with Gravitational Waves

### v1

@@ I. AD scalars: background dynamics and asymmetry generation
We now present some numerical examples of the asymmetry. For $\epsilon_{\Phi}=10^{-20}$ the observed $n_{B}/s\sim 10^{-10}$ can be realized with $T_{d}\sim 2.5\times 10^{7}$ GeV when $m_{\Phi}=10\,\mathrm{GeV}$. As we will discuss, transfer of the asymmetry to the baryon sector may involve significant washout (model-dependent), requiring a larger initial $\Phi$ asymmetry to compensate. Therefore we also consider the parameter points that lead to $n_{\Phi}/s\gg 10^{-10}$, for example, with the same $\epsilon_{\Phi}$, but $T_{d}\sim 1$ GeV, one obtains $n_{\Phi}/s\sim 4\times 10^{-5}$. A more detailed parameter scan is given in Section .3 of the Supplemental Material and Fig. S1, demonstrating the possibility of producing a large initial asymmetry to compensate for potential (model-dependent) washout effects.
II. Transfer to SM baryon asymmetry.
@@
The asymmetry in the $\Phi$ field generated during its oscillation needs to eventually be transferred to the SM baryon asymmetry via its subsequent decay to baryons or leptons (further assisted by electroweak sphalerons in the latter case). This can generally be realized by a B- or L-violating coupling involving $\Phi$ in terms of effective higher dimensional operators, such as $\Phi QQQL/\Lambda^{3}$ Allahverdi and Mazumdar (2012), $\Phi UDDUDD/\Lambda^{6}$ and $\Phi LLELLE/\Lambda^{6}$. Whether the related constraints, for instance, from $n-\bar{n}$ oscillation, can be satisfied depends on the specifics of the UV completion, including flavor structure in the couplings. As a proof of principle, in the following, we outline examples of renormalizable UV complete models realizing the $\Phi$ to baryon asymmetry transfer and summarize the conditions on viable model parameters. Further details of the analysis are given in Section .4 of the Supplemental Material. The core results of this work are independent of the details of the asymmetry transfer mechanism.
@@
where $N_{i}$’s are right-handed (RH) Majorana neutrinos (with $i=1,2,3$ as a typical choice), $H$ is SM Higgs, $L_{j}$ are SM lepton doublets. This model potentially relates to neutrino mass generation, which nevertheless is not imposed as a requirement here. We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes enabled by the $NHL$ coupling, as considered in the literature of leptogenesis Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Hambye and Teresi (2016); Flood et al. (2022)  [^4] Unlike in leptogenesis, here we do not require additional CP violation or out-of-equilibrium condition in neutrino sector to generate the asymmetry, as the asymmetry is already generated in $\Phi$..
@@
A simple renormalizable model, in this case, includes:
@@
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. In the viable parameter region, the decay proceeds with $\Phi\rightarrow\chi\chi$, followed by $\chi\rightarrow udd$.
Both models are subject to cosmological conditions which impose lower limits on $m_{\Phi}$ and $m_{\chi}$ or $m_{N}$. For $\Phi$ to decay to baryons (before BBN), it is subject to kinematic conditions: $m_{\Phi}\gtrsim 2$ GeV, and $m_{\chi}\gtrsim 1$ GeV. For $\Phi$ asymmetry to be transferred to the SM leptons via $N$ before the EW phase transition, $m_{\Phi}>2m_{N}\gtrsim O(0.1)$ GeV Klarić et al. (2021a). In addition, for both models, the potential asymmetry washout effects should be either suppressed or can be compensated by a larger initial $\Phi$ asymmetry produced during its oscillation (which has been shown viable in the last section). Furthermore, relevant laboratory experiments at energy and intensity frontiers searching for B-violation, color-charged new particles, sterile neutrinos, etc., may provide additional constraints on model parameters. In Section .4 of the Supplemental Material, we elaborate on relevant constraints and demonstrate examples of viable parameter regions. On the other hand, as we will see in our GW results, an upper limit in mass of $m_{\Phi}\lesssim O(10)$ GeV corresponds to the currently detectable GW peak frequency range, which is below $\sim O(100)$ Hz. This, combined with the aforementioned cosmological lower limit of $m_{\Phi}\gtrsim O(0.1-1)$ GeV, demonstrates that intriguingly the parameter range of our interest points to new physics at the energy scale of $O(0.1-10)$ GeV, indicating a new potential complementarity between GW detection and laboratory searches across particle physics frontiers.
Parametric Resonance and GW Signals. We have seen that the generation of the baryon asymmetry is determined by the classical motion of $\Phi$ as an averaged background field. In addition to this classical motion of $\Phi$, we must consider the corresponding fluctuations around it, which leads to parametric resonance that may source GWs. We proceed by decomposing $\phi_{R,I}=\bar{\phi}_{R,I}(t)+\delta\phi_{R,I}(x,t)$, where $\bar{\phi}_{R,I}$ describe the classical background motion and $\delta\phi_{R,I}(x,t)$ correspond to the fluctuations, which have a quantum origin. The evolution of background fields $\bar{\phi}_{R,I}$ follows Eqs. (2), (3), whereas we expand the fluctuations in Fourier modes as
@@
As the background fields oscillate around the minimum of the potential, the fluctuations in momentum-space evolve according to
@@
$+\lambda_{\phi}(3\bar{\phi}_{R,I}^{2}+\bar{\phi}_{I,R}^{2})\delta\phi_{R,I}+2\bar{\phi}_{I}\bar{\phi}_{R}\delta\phi_{I,R}=0$ (9)
@@
In this case, we can rotate our field basis and define fields in the radial direction (along the background motion) ϕsubscriptitalic-$\phi_{\parallel}$ and perpendicular to it (orthogonal direction) as ϕsubscriptitalic-$\phi_{\perp}$. By construction, $\bar{\phi}_{\perp}=0$, and the fluctuation equations for ϕsubscriptitalic-$\phi_{\parallel}$ and ϕsubscriptitalic-$\phi_{\perp}$ decouple, namely,
$\ddot{\delta\phi}_{\parallel,\perp}+3H\dot{\delta\phi}_{\parallel,\perp}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+q_{\parallel,\perp}\lambda_{\Phi}\bar{\phi}_{\parallel,\perp}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (10)
where $q_{\parallel,\perp}=\{3,1\}$. As the background field oscillates quasi-periodically around the origin, fluctuations can be exponentially amplified for certain ranges of wavenumbers. Through the exponential amplification, the occupation number of these modes grows rapidly, and they can thus be treated classically Polarski and Starobinsky (1992); Kofman et al. (1997). This phenomenon is referred to as “preheating” Greene et al. (1997) in the context of primordial particle production following inflation, while it also can occur for oscillating spectator scalar fields in general (see, for instance, our related work Cui et al. (2024)).
To quantify parametric resonance, we start by defining a dimensionless field amplitude $\tilde{\phi}=\phi/\phi_{\mathrm{in}}$ and a corresponding conformal scaling for time as $\mathrm{d}\eta=\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}\mathrm{d}t/a$ with $\phi_{\mathrm{in}}$ being the initial amplitude of the scalar field $\phi_{\mathrm{in}}=|\Phi|_{\mathrm{in}}$. The dimensionless fluctuation equation becomes (derivatives are over $\eta$)
${\delta\phi}_{\parallel,\perp}^{\prime\prime}+2\tilde{H}{\delta\phi}_{\parallel,\perp}^{\prime}+\left({\tilde{k}^{2}}+a^{2}\mu^{2}+q_{\parallel,\perp}a^{2}\tilde{\phi}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (11)
@@
[…]
The growth of fluctuations and their back-reaction break the homogeneous oscillation of $\Phi$, leading to the inhomogeneous structure of scalar fields with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{a}\nabla_{j}\phi^{a}$ (where the repeated index `$a$’ indicates the number of the scalar components and a summation is implied). These time-dependent field inhomogeneities source (sub-Hubble) gravitational waves whose frequency depends on the characteristic energy scale involved $H_{c}$. The peak amplitude and frequency of GW spectra generated from sources peaked at some characteristic scale $k_{c}$ in momentum-space can be estimated as Giblin and Thrane (2014)
$f_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.7\times 10^{10}\sqrt{\frac{H_{c}}{M_{\mathrm{P}}}}\frac{k_{c}}{H_{c}}\,{\mathrm{Hz}}\,,$ (12)
$\Omega_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.3\times 10^{-4}\,\alpha^{2}\,\beta\,w\left(\frac{k_{c}}{\sigma}\right)\left(\frac{H_{c}}{k_{c}}\right)^{2}.$ (13)
where $\alpha$ is a measure of the fraction of the energy in the GW source here the scalar field inhomogeneities and this $\alpha$ coincides with the $\alpha$ we defined below Eq. 5 relative to the total energy density of the Universe, and $\beta$ describes the degree of anisotropy of the source which must be determined from simulations. In typical scenarios $\beta=\mathcal{O}(0.01-0.1)$. The equation of the state of the Universe at the time of the production of GWs for a radiation-dominated Universe is $w=1/3$ and $\sigma$ is the width of the source in momentum-space, which can be taken as $\sigma\sim k_{c}$ for peaked sources (see Ref. Giblin and Thrane (2014) for the derivation and e.g. Refs. Cui and Sfakianakis (2023); Cui et al. (2024) for further applications of Eqs (12) and (13)).
The dynamics of the generated GWs are governed by the linearized equation for the transverse-traceless tensor perturbation sourced by the above anisotropic stress
$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{\nabla^{2}h_{ij}}{a^{2}}=\frac{2}{a^{2}M_{\mathrm{P}}^{2}}\Pi_{ij}^{TT},$ (14)
The quantity of interest is the energy density power spectrum of the GWs normalized by the critical energy density of the Universe today. For GWs produced in a radiation-dominated era, the spectrum of the energy density of GWs (per logarithmic momentum interval) observable today is Easther and Lim (2006); Easther et al. (2007, 2008); Dufaux et al. (2007); Cui et al. (2024); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al. (2008):
@@
where the critical density of the Universe is $\rho_{\mathrm{crit}}=3M_{\mathrm{P}}^{2}H^{2}$ and $\Omega_{\mathrm{rad},0}h^{2}=h^{2}\rho_{\mathrm{rad},0}/\rho_{\mathrm{crit}}=4.3\times 10^{-5}$. In particular, we employ the publicly available code Cosmolattice Figueroa et al. (2021, 2023) in a three-dimensional lattice, in order to numerically solve the metric perturbation equations in Eq. (14) along with the equations for the scalar system in an expanding universe.
@@
The benchmark examples of observed spectra today from our model simulation are shown in Fig 1. For the initially $\Phi^{4}$ dominant system, as in our case, the field starts rolling/oscillating when $H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$. This combination of $\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ also determines the frequency of the observed GWs, according to Eq. 12. We consider the RD universe with the scalar field system contributing to a fraction of the total energy $\alpha$ and $\alpha=0.3$ is assumed in Fig 1. According to Eq. 13, $\Omega^{\mathrm{peak}}_{\mathrm{GW}}\propto\alpha^{2}\propto\phi^{8}_{\mathrm{in}}$, and we found that a sizable $\alpha\sim O(0.1)$ which optimizes GW amplitude requires $\phi_{\mathrm{in}}\sim M_{\mathrm{P}}$. As we have seen from the discussion on the resonance structure of the model, the parameters $\mu_{\pm}$ (or simply $\mu$ when $A=0$) determine the nature and efficiency of resonance Greene et al. (1997). We take $\mu=10^{-5}$ as an example for stable resonance, which fixes $\lambda_{\Phi}\phi_{\mathrm{in}}^{2}$. with the optimal choice of $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$, we find $\lambda_{\Phi}\sim 10^{-30}$ for the mass range of interest. As discussed earlier, $m_{\Phi}\gtrsim O(0.1)$ GeV is required for $\Phi$ to successfully transfer the asymmetry to SM baryons. Interestingly, in the viable range of $m_{\Phi}\sim O(0.1)-O(10)$ GeV, the resulting peak GW frequency lies well within the range of ET/CE and with sufficient amplitudes, making signal detection promising for the near future. A higher range of $m_{\Phi}$ is compatible with baryogenesis but requires futuristic higher frequency GW detectors Aggarwal et al. (2021) to capture the peak frequency range of the signal. Due to the finite box size and lattice spacing, low-frequency modes well off the peak cannot be resolved in our simulation, which is a common issue in the lattice simulation literature Easther et al. (2007). Therefore, to extend the spectral prediction to lower frequencies, we apply the standard assumption that at low wave numbers, the GW spectrum grows as $k^{3}$, as derived from causality arguments Easther et al. (2007); Caprini et al. (2009); Hook et al. (2021). This is shown in dashed lines in Fig. 1. Depending on $m_{\Phi}$, the low-$k$ tail may be detectable by DECIGO and BBO, providing the prospect for a correlated detection across multiple frequency bands, probing the peak and tail of the spectrum simultaneously.
@@
Figure 1: We show the GW spectrum originating from the AD model considered here for three benchmark masses of $\Phi$. The potential parameters $\lambda_{\Phi}=10^{-30}$, $A=0$ are chosen to ensure efficient parametric resonance (see main text for details). The fraction of initial scalar field energy density over the total energy density of the Universe is $\alpha=30\%$. The dashed lines at lower frequencies are extrapolated according to the causal super-horizon $k^{3}$ scaling.
Finally, note that since $\Phi$ is coupled to SM fermions in order to transfer the asymmetry to baryons, the effect of parametric excitation of these fermionic states should be evaluated. Nevertheless, in the context of preheating, it was found that during scalar field oscillations, the energy transfer to bosons is typically faster and larger than to fermions due to Pauli blocking Greene and Kofman (1999); Giudice et al. (1999); Greene and Kofman (2000); Sorbo (2001); Adshead and Sfakianakis (2015). Therefore, we expect our resonance analysis to remain robust in the presence of fermion couplings.
@@
Conclusion and Outlook. In this Letter, we first demonstrate that the Affleck-Dine mechanism for baryogenesis can be realized by a complex light scalar field $\Phi$ with a mass $m_{\Phi}$ well below the electroweak scale without invoking supersymmetry. In particular, we use a convex, renormalizable, polynomial potential (which does not lead to the formation of Q-balls generally present in SUSY AD models). Such a scalar field acts as a spectator field during inflation while acquiring a large field value at the end of inflation. The subsequent rolling down and oscillation of the scalar field occur long after inflation, during the radiation-dominated era, triggering the generation of baryon asymmetry. Parametric resonance and fragmentation generally arise during the oscillation of the scalar field, naturally leading to GW production. We demonstrate that the theory may lead to successful baryogenesis while sourcing detectable GW signals with a peak frequency of O(10-100) Hz, within reach of experiments such as ET and CE. Furthermore, the characteristic new physics scale, characterized by $m_{\Phi}$, intriguingly lies in the range of ${\cal O}(0.1\mathrm{-}10)$ GeV, while the transfer of the $\Phi$ asymmetry to the SM B- or L-asymmetry requires interactions between the new physics sector and the SM states. Hence, from this well-motivated scenario, a new, natural complementarity arises between SGWB detection and laboratory searches for new particle physics across the energy, intensity, and neutrino frontiers. The specifics of the complementary laboratory signal depend on the details of the asymmetry transfer mechanism. Examples include: searches for ${\cal O}(0.1\mathrm{-}10)$ GeV mass sterile neutrino with DUNE, SHiP, FASER, etc. and exotic multi-jet signals at the LHC (see Section .4 of the Supplemental Material). This work thus sheds new light on probing early Universe baryogenesis through its GW imprint, potentially complemented by associated laboratory probes, while presenting a well-motivated new physics source for SGWB.

### v2

@@ Abstract
We now present some numerical examples of the asymmetry. For $\epsilon_{\Phi}=10^{-20}$ the observed $n_{B}/s\sim 10^{-10}$ can be realized with $T_{d}\sim 2.5\times 10^{7}$ GeV when $m_{\Phi}=10\,\mathrm{GeV}$. As we will discuss, transferring the asymmetry to the baryon sector may involve a significant (model-dependent) washout, requiring a larger initial $\Phi$ asymmetry to compensate. Therefore, we also consider parameter points that lead to $n_{\Phi}/s\gg 10^{-10}$, for example, the same $\epsilon_{\Phi}$ with $T_{d}\sim 1$ GeV gives $n_{\Phi}/s\sim 4\times 10^{-5}$. A more detailed parameter scan is provided in Section C of Supplemental Material, demonstrating the potential to produce a large initial asymmetry that compensates for possible washout effects.
II. Transfer to the SM baryon asymmetry .
@@
The asymmetry in the $\Phi$ field generated during its oscillation needs to eventually be transferred to the SM baryon asymmetry via its subsequent decay to baryons or leptons (further assisted by electroweak sphalerons in the latter case). This can generally be realized by a B- or L-violating coupling involving $\Phi$ in terms of effective higher-dimensional operators, such as $\Phi QQQL/\Lambda^{3}$ Allahverdi and Mazumdar (2012), $\Phi UDDUDD/\Lambda^{6}$, and $\Phi LLELLE/\Lambda^{6}$. Whether the related constraints, for instance, from $n-\bar{n}$ oscillation, can be satisfied depends on the specifics of the UV completion, including flavor structure in the couplings. As a proof of principle, we outline examples of renormalizable, UV-complete models that realize the $\Phi$ to baryon asymmetry transfer in the following, and summarize the conditions on viable model parameters. Further details are given in Section D of Supplemental Material. The core results of this work are independent of the details of the asymmetry transfer mechanism.
@@
where $N_{i}$’s are right-handed Majorana neutrinos (with $i=1,2,3$ as a typical choice), $H$ is SM Higgs, and $L_{j}$ are SM lepton doublets. This model potentially relates to neutrino mass generation, which, nevertheless, is not imposed as a requirement here. We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes enabled by the $NHL$ coupling, as considered in the literature of leptogenesis Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Hambye and Teresi (2016); Flood et al. (2022)  [^6] Unlike in leptogenesis, here we do not require additional CP violation or out-of-equilibrium condition in neutrino sector to generate the asymmetry, as the asymmetry is already generated in $\Phi$..
@@
A simple renormalizable model, in this case, includes
@@
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices i,j=1,2,3i,~j=1,2,3 in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. In the viable parameter region, the decay proceeds with $\Phi\rightarrow\chi\chi$, followed by $\chi\rightarrow udd$.
Both models are subject to cosmological conditions that impose lower limits on $m_{\Phi}$, $m_{\chi}$, or $m_{N}$. For $\Phi$ to decay to baryons (before BBN), it is subject to kinematic conditions: $m_{\Phi}\gtrsim 2$ GeV, and $m_{\chi}\gtrsim 1$ GeV. For $\Phi$ asymmetry to be transferred to SM leptons via $N$ before the EW phase transition requires $m_{\Phi}>2m_{N}\gtrsim O(0.1)$ GeV Klarić et al. (2021a). In addition, for both models, the potential asymmetry washout effects should be either suppressed or compensated by a larger initial $\Phi$ asymmetry produced during its oscillation (which has been shown to be viable in the last section). Furthermore, relevant laboratory experiments at the energy and intensity frontiers, searching for B-violation, color-charged new particles, and sterile neutrinos, among others, may provide additional constraints on model parameters. In Section D of Supplemental Material, we elaborate on relevant constraints and demonstrate examples of viable parameter regions.
Parametric Resonance and GW Signals. We have seen that the generation of the baryon asymmetry is determined by the classical motion of $\Phi$ as an averaged background field. In addition to this classical motion of $\Phi$, the corresponding fluctuations around it can lead to parametric resonance and source GWs. We decompose $\phi_{R,I}=\bar{\phi}_{R,I}(t)+\delta\phi_{R,I}(x,t)$, where $\bar{\phi}_{R,I}$ describe the classical background motion and $\delta\phi_{R,I}(x,t)$ correspond to the fluctuations, which have a quantum origin. The evolution of background fields $\bar{\phi}_{R,I}$ follows Eqs. (2) and (3), whereas we expand the fluctuations in Fourier modes as
@@
As the background fields oscillate around the minimum of the potential, the fluctuations in momentum space evolve according to
@@
$\quad+\lambda_{\phi}(3\bar{\phi}_{R,I}^{2}+\bar{\phi}_{I,R}^{2})\delta\phi_{R,I}+2\bar{\phi}_{I}\bar{\phi}_{R}\delta\phi_{I,R}=0.$ (12)
@@
In this case, we can rotate our field basis and define fields in the radial direction (along the background motion) $\phi_{\parallel}$ and perpendicular to it (orthogonal direction) as $\phi_{\perp}$. By construction, $\bar{\phi}_{\perp}=0$, and the fluctuation equations for $\phi_{\parallel}$ and $\phi_{\perp}$ decouple, namely,
$\ddot{\delta\phi}_{\parallel,\perp}+3H\dot{\delta\phi}_{\parallel,\perp}+{k^{2}\over a^{2}}{\delta\phi}_{\parallel,\perp}+m_{\Phi}^{2}{\delta\phi}_{\parallel,\perp}$
$+q_{\parallel,\perp}\lambda_{\Phi}\bar{\phi}_{\parallel,\perp}^{2}\delta\phi_{\parallel,\perp}=0,$ (13)
where $q_{\parallel,\perp}=\{3,1\}$. Quasiperiodic background-field oscillations can exponentially amplify fluctuations within certain wave-number ranges. This rapid growth of mode occupation numbers allows these modes to be treated classically Polarski and Starobinsky (1992); Kofman et al. (1997). This phenomenon, originally studied for exponential particle production (“preheating”) after inflation Greene et al. (1997), can also occur for any oscillating scalar (see, for instance, Cui et al. (2024)).
To quantify parametric resonance, we define a dimensionless field amplitude $\tilde{\phi}=\phi/\phi_{\mathrm{in}}$ and a corresponding conformal scaling for time as $\mathrm{d}\eta=\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}\mathrm{d}t/a$ with $\phi_{\mathrm{in}}$ being the initial amplitude of the scalar field $\phi_{\mathrm{in}}=\left|\Phi_{\mathrm{in}}\right|$. The dimensionless fluctuation equation becomes ($\delta\phi^{\prime}\equiv d\delta\phi/d\eta$):
${\delta\phi}_{\parallel,\perp}^{\prime\prime}+2\tilde{H}{\delta\phi}_{\parallel,\perp}^{\prime}+\left({\tilde{k}^{2}}+a^{2}\mu^{2}+q_{\parallel,\perp}a^{2}\tilde{\phi}^{2}\right)\delta\phi_{\parallel,\perp}=0,$ (14)
@@
There are two essential resonance parameters: $\mu^{2}=\frac{m_{\Phi}^{2}}{\lambda_{\Phi}\phi_{\mathrm{in}}^{2}}$ and $q_{\parallel,\perp}$. As shown in Greene et al. (1997), a large $\mu^{2}$ can suppress the resonance and determine if the system displays stable resonance, stochastic resonance, or no decay. We choose to focus on $\mu\leq 10^{-5}$, such that the system is always in a stable resonance regime. For such small $\mu^{2}$, the system is similar to the well-studied case of a quartic field during preheating, whereas $q_{\parallel,\perp}$ is fixed in our case due to the rotational symmetry of the potential for $A=0$. Section E of Supplemental Material discusses in detail the effect of $q_{\parallel,\perp}$ on the growth of $\delta\phi_{\parallel,\perp}$ and shows that for sufficiently small values of the asymmetry parameter $A\lesssim 0.005m_{\Phi}^{2}$, the parametric resonance dynamics is indistinguishable from the case $A=0$, and thus the GW amplitude remains insensitive to the value of $A$. Fortunately, the parameter range of $A$ that produces sufficient baryon asymmetry, as considered in our benchmarks, is well within this regime. On a technical note, we use the parametrization of real and imaginary components of $\Phi$ ($\phi_{R}$, $\phi_{I}$) when computing the baryon asymmetry, whereas we are using the radial direction and its orthogonal ($\phi_{\parallel}$, $\phi_{\perp}$) when computing parametric resonance for $A\ll m_{\Phi}^{2}$. This is due to the fact that the baryon asymmetry scales with $A$ and thus the difference between the $\phi_{I}$ and $\phi_{R}$ directions is important, whereas parametric resonance and GW production are blind to $A\ll m_{\Phi}^{2}$, making radial coordinates more suitable. Interesting deviations from the $A=0$ limit are discussed in Section E of Supplemental Material. For $A\gtrsim 0.05m_{\Phi}^{2}$ the GW spectrum starts being sensitive to the exact value of $A$ as well as the initial angle $\theta_{\mathrm{in}}$. This provides intriguing possibilities for distinguishing between these parameters by dedicated analysis of the GW spectral shape.
The growth of fluctuations and their backreaction breaks the homogeneous oscillation of $\Phi$, leading to inhomogeneities with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{a}\nabla_{j}\phi^{a}$ (where the repeated index a`a’ indicates the number of the scalar components and a summation is implied). The dynamics of the generated GWs are governed by the linearized equation for the transverse-traceless tensor perturbation sourced by the above anisotropic stress
$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{\nabla^{2}h_{ij}}{a^{2}}=\frac{2}{a^{2}M_{\mathrm{P}}^{2}}\Pi_{ij}^{TT}\,.$ (15)
The quantity of interest is the energy density power spectrum of the GWs normalized by the critical energy density of the Universe today. For GWs produced in a radiation-dominated era, the spectrum of the energy density of GWs (per logarithmic momentum interval) observable today is Easther and Lim (2006); Easther et al. (2007, 2008); Dufaux et al. (2007); Cui et al. (2024); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al. (2008)
@@
where the critical density of the Universe is $\rho_{\mathrm{crit}}=3M_{\mathrm{P}}^{2}H^{2}$ and $\Omega_{\mathrm{rad},0}h^{2}=h^{2}\rho_{\mathrm{rad},0}/\rho_{\mathrm{crit}}\simeq 4.3\times 10^{-5}$. In particular, we employ the publicly available code Cosmolattice Figueroa et al. (2021, 2023) in a three-dimensional lattice, in order to numerically solve the metric perturbation equations in Eq. (15) along with the equations for the scalar system in an expanding universe.
@@
The spectra for three benchmark examples are shown in Fig 1. For a $\Phi^{4}$-dominated system, as in our case initially, the field starts rolling/oscillating when $H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$. The combination $\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ also determines the frequency of the observed GWs. We consider the RD Universe with the scalar field system contributing a fraction of the total energy $\alpha$; Fig 1 corresponds to $\alpha=0.3$. For GWs generated by sources peaked around a characteristic scale $k_{c}$ in the momentum space Giblin and Thrane (2014), $\Omega^{\mathrm{peak}}_{\mathrm{GW}}\propto\alpha^{2}\propto\phi^{8}_{\mathrm{in}}$, and we found that a sizable $\alpha\sim O(0.1)$ which optimizes GW amplitude requires $\phi_{\mathrm{in}}\sim M_{\mathrm{P}}$ (see also Cui et al. (2024)). Furthermore, the observed frequency today is $f_{\mathrm{GW}}^{\mathrm{peak}}\simeq 2.7\times 10^{10}k_{c}/\sqrt{H_{c}M_{\mathrm{Pl}}}$ Hz, where $k_{c}\sim H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ (see Section E of Supplemental Material for details), leading to $f_{\mathrm{GW}}^{\mathrm{peak}}\sim 10^{10}\lambda_{\Phi}^{1/4}$ for $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$. As we have seen from the discussion on the resonance structure of the model, the parameters $\mu_{\pm}$ (or simply $\mu$ when $A=0$) determine the nature and efficiency of resonance Greene et al. (1997). We found that stable resonance occurs for $\mu\equiv{m_{\Phi}}/({\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}})\lesssim 10^{-5}$. We choose $\mu=10^{-5}$, which fixes the product $\lambda_{\Phi}\phi_{\mathrm{in}}^{2}$. With an optimal initial value of $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$, we find that $\lambda_{\Phi}\simeq 10^{-35},10^{-33},10^{-31}$ for $m_{\Phi}=0.1,1,10$ GeV; see Fig. 1. As discussed earlier, $m_{\Phi}\gtrsim O(0.1)$ GeV is required for $\Phi$ to successfully transfer the asymmetry to SM baryons. Interestingly, in the viable range of $m_{\Phi}\sim O(0.1)-O(10)$ GeV, the resulting peak GW frequency lies well within the range of ET/CE and with sufficient amplitudes, making signal detection promising for the near future. A higher range of $m_{\Phi}$ is compatible with baryogenesis but requires futuristic, higher-frequency GW detectors Aggarwal et al. (2021) to capture the peak frequency range of the signal. Due to the finite box size and lattice spacing, low-frequency modes well off the peak cannot be resolved in our simulation, which is a common issue in the lattice simulation literature Easther et al. (2007). Therefore, to extend the spectral prediction to lower frequencies, we apply the standard assumption that at low wave numbers, the GW spectrum scales as $k^{3}$, as derived from causality arguments Easther et al. (2007); Caprini et al. (2009); Cai et al. (2020); Hook et al. (2021). This is shown in dashed lines in Fig. 1. Depending on $m_{\Phi}$, the low-$k$ tail may be detectable by DECIGO and BBO, providing the prospect for correlated detection across multiple frequency bands, which simultaneously probes the peak and tail of the spectrum. Before concluding, let us note that the simulated signal shown in Fig. 1 is ${\cal O}(100)$ times larger than the lowest point of the CE sensitivity curve. Since $\Omega_{GW}h^{2}\propto\alpha^{2}$ Giblin and Thrane (2014), we can reduce $\alpha$, the initial fraction of the energy density of the scalar field with respect to the total energy density in the Universe, down to $\alpha\sim O(1\%)$ and still yield a GW signal within reach of CE.
@@
Figure 1: We show the GW spectrum originating from the AD model considered here for three benchmark masses of $\Phi$. The potential parameters $\lambda_{\Phi}\sim 10^{10}m_{\Phi}^{2}/M_{\mathrm{Pl}}^{2}$, $A=0$ are chosen to ensure efficient parametric resonance (see the main text for details). The fraction of initial scalar field energy density over the total energy density of the Universe is $\alpha=30\%$. The dashed lines at lower frequencies are extrapolated according to the causal superhorizon $k^{3}$ scaling.
@@
Conclusion and Outlook. In this Letter, we first demonstrate that the Affleck-Dine mechanism for baryogenesis can be realized by a complex light scalar field $\Phi$ with a mass $m_{\Phi}$ well below the electroweak scale without invoking supersymmetry. In particular, we use a convex, renormalizable, polynomial potential (which does not lead to the formation of Q-balls, generally present in SUSY AD models). Such a scalar field acts as a spectator field during inflation while acquiring a large field value at the end of inflation. The subsequent rolling down and oscillation of the scalar field occur long after inflation, during the radiation-dominated era, triggering the generation of baryon asymmetry. Parametric resonance and fragmentation generally arise during the oscillation of the scalar field, naturally leading to GW production. We demonstrate that the models considered can lead to successful baryogenesis while sourcing detectable GW signals with a peak frequency of O(10-100) Hz, within reach of experiments such as the ET and CE. Furthermore, the characteristic new physics scale, characterized by $m_{\Phi}$, intriguingly lies in the range of ${\cal O}(0.1\mathrm{-}10)$ GeV, while the transfer of the $\Phi$ asymmetry to the SM B- or L-asymmetry requires interactions between the new physics sector and the SM states. Hence, from this well-motivated scenario, a new, natural complementarity arises between SGWB detection and laboratory searches for new particle physics across the energy, intensity, and neutrino frontiers. The specifics of the complementary laboratory signal depend on the details of the asymmetry transfer mechanism. Examples include searches for ${\cal O}(0.1\mathrm{-}10)$ GeV mass sterile neutrino with DUNE, SHiP, FASER, etc., and exotic multijet signals at the LHC (see Supplemental Material, Section D). This work thus sheds new light on probing early Universe baryogenesis through its GW imprint, potentially complemented by associated laboratory probes, while presenting a well-motivated new physics source for SGWB.